# A Survey of Conversational Search

### v1

@@ 4.4 Interpretation in Conversational Dense Retrieval
Although conversational dense retrieval has proven effective for conversational search, a significant limitation of these approaches is their lack of interpretability, which impedes an intuitive understanding of model behaviors and hinders targeted improvements. To this end, the existing studies [243] attempt to unravel the behavior characteristics of various conversational dense retrieval models. For example, Mao et al. [131] propose LeCoRE, which enhances interpretability by creating a denoised and sparse lexical representation of the conversational session. Specifically, LeCoRE extends the SPLADE model [240], a lexical-based model to further transform latent representation into lexical representation in the vocabulary space, which provides a clear and understandable representation related to the search results. Besides, Christmann et al. [132] increase the model interpretability by combining information from multiple sources with user-comprehensible explanations for search results. Further, Cheng et al. propose CONVINV [133] to transform the opaque, high-dimensional session embeddings generated by conversational dense retrieval models into explicit, interpretable text that can be easily understood by humans. The transformation is achieved by a model trained to invert the session embeddings back into a textual form that is not only coherent and clear but also maintains the retrieval performance of the original embeddings. Instead of explicitly conducting query reformulation for conversational search, more detailed investigations of the interpretability in dense retrieval are expected, which can avoid unfair or misleading search results without the opportunity for timely correction.
@@ 4.5 Re-ranking in Conversational Search
In the context of re-ranking in conversational search, Kongyoung et al. [134] design monoQA to train a model with multi-task learning simultaneously for re-ranking retrieved passages and extracting answers. By integrating re-ranking and reading into a unified model, monoQA improves overall retrieval precision and answer accuracy. Ju et al. [28] attempt to improve re-ranking accuracy by employing a multi-view pseudo-labeling approach that integrates diverse contextual information. It utilizes a view-ensemble method to generate high-quality pseudo-labeled data for fine-tuning a conversational passage re-ranker. Hai et al. [118] design a re-ranking process after the conversational retrieval stage, which is enriched with keywords identified by the first-stage retriever. These keywords are expected to capture the helpful contextual information and re-rank the documents with useful conversational history. Owoicho et al. [135] investigate conversational search systems by providing feedback on system responses and answering clarifying questions, based on an initial description of the user’s information need. This work discusses the potential advancements in conversational search by developing more sophisticated feedback processing modules, e.g., conversational context-originated re-ranker.
@@ 4.6 Limitations and Discussion
Standard of Effective Pipeline. Though various pipelines with different mechanisms for conversational retrieval have been explored in the literature, how to establish a unified standard to construct and compare effective pipelines is still an open question, e.g., how to compare query reformulation approaches with conversational dense retrieval models. Besides, how to incorporate the multi-turn information, e.g., the historical retrieved results into the current turn retrieval is also under explored, as well as the necessity of deploying a re-ranker in the retrieval pipeline. These problems require the community to conduct further investigation on how to build a conversational retrieval pipeline with the associated historical context under a comparable standard.
@@
Explainability. Though some existing works begin to pay attention to the explainability of conversational dense retrievers, the investigation of the interpretation studies is still limited. Detailed and thorough studies on conversational (dense) retrieval models can help the community understand the behavior of both the user and the retriever, which is essential for an interactive system. One potential solution is to conduct a comprehensive performance analysis that explores how different aspects of the conversation (e.g., topic switch, expansion utilization) affect retrieval effectiveness and user modeling.
@@ 5 Generation in Conversational Search
The development of generative models facilitates the evolution of conversational search systems that might enrich the user experience by producing contextually aware and personalized responses [2]. As the chat mode, i.e., multi-turn conversation, becomes a common practice when leveraging LLMs for content generation, the design of the new paradigms of the conversational search system by integrating the generation module brings much more attention, however, which is not well-studied in the literature.
@@
The existing studies mainly focus on single-turn retrieval-augmented generation (RAG), which only treats the single stand-alone query or converts the multi-turn context-dependent query by reformulation technique without leveraging historical information for RAG. In terms of the single-turn RAG paradigm, the generation component is pivotal for crafting responses that are both informed by external knowledge and coherent with the parametric ones. The retrieved documents usually served as external knowledge, and the generator model, usually the LLM in existing studies, processes them with two key methodologies [245]: context curation [246, 247, 248, 249] and generator fine-tuning [250, 251], which try to adapt the consistency between IR and RAG [252, 253] and satisfy the specific requirements of the domain or task, respectively.
@@
Regarding the generation for conversational search, the single-turn RAG systems overlook the historical information, e.g., historical search results, the context history, turn dependency, etc. Since the conversational search systems must account for the ongoing dialogue and the evolving context of the conversation to produce responses that are not only relevant but also consistent with the user’s previous interactions, historical information might be essential to generate accurate, reliable, and personalized responses. Thus, it is unclear whether the additional conversational context process can further improve the generation results, even though some mechanisms with similar goals have already been applied in the retrieval procedure. Considering few studies are focusing on the concept of generation in conversational search, in this section, we aim to provide the potential research questions, guidance for unclear points, and corresponding solutions with different aspects to facilitate future research in this field. An overview of the conversational generation system is illustrated in Figure 7.
@@
Figure 7: Overview of the conversational response generation system. To achieve more user-centric and reliable responses, potential directions that are worth further exploring include historical context utilization, context dependency modeling, and knowledge attribution.
@@ 5.3 Conversational Knowledge Attribution
Although these methods enhance knowledge attribution and interpretability from different perspectives, they are not designed for conversational scenarios. Thus, it is unclear whether they can be applied to the conversational generation. For example, could these methods also be attributed to the historical evidence retrieved when injecting context information from previous turns? The potential exploration and improvement could begin with the following aspects.
@@ 6 Domain-specific and User-centric Conversational Search
The rapid development of conversational search systems facilitates various applications in the general domain and across diverse specific domains, including medical [261], finance [262], law [51], e-commercial [145], etc. The domain-specific systems might store the domain-specific knowledge in different forms, e.g., tables, graphs, or multi-modal, which require the search systems to gather various information resources via reasoning skills. Aside from specific domain requirements, the system should cover the intended utilization of the conversational search model by real users. Thus, the designed system should also consider a user-centric perspective [263] in both dataset construction and evaluation criteria rather than only focusing on predefined search abilities. Besides, from the personalized perspective, since the backgrounds of users are different, the search results from the conversational systems should yield different results for the same query according to the user profiles based on identifying the relevant information in the user profiles for each query turn [10]. In this section, we present the pioneering conversational search studies conducted in various specific domains and user-centric scenarios, and we conclude by offering a comprehensive discussion of the emerging limitations and challenges.
@@ 6.1 Domain-specific Conversational Search
Earlier studies [264, 265, 266] attempt to solve domain-specific issues via multi-domain/task learning based on task-originated conversation system. The key idea is to search cross-domain information for sharing by reasoning over the conversational history. Though these systems can somehow process various types of information across domains, they cannot achieve satisfactory performance when targeting each specific domain. Thus, various domain-specific models are proposed for each domain.
@@
The common practice of establishing a conversational search system in the medical domain is to extract relevant medical concepts from retrieved information within a conversational interface for response generation [33], which is illustrated in figure 8 (a). The variability in how medical concepts are expressed (i.e., the same concept might be written in numerous non-canonical forms) poses challenges for accurately searching texts with relevant concepts and for modeling conversational context, particularly with co-reference and omission [34].
@@
With the emergence of LLMs, the medical domain conversational search systems paradigm encounters new opportunities and challenges [267, 261]. The available medical corpus [35, 36] enables continual training of the existing LLMs to enhance the superb capability on the associated domain. For example, Toma et al. [37] propose an open-source expert-level medical language model by encoding the knowledge based on conversational context. Bao et al. [38] construct high-quality supervised fine-tuning datasets based on dialogue and knowledge graphs to bridge LLMs with real-world medical consultation. However, the generative LLMs remain an unyielding concern in returning accurate responses, especially when the stored knowledge is out-to-date. To mitigate such issues, Pal et al. [39] propose medical domain hallucination detection approaches for LLMs to reduce inaccurate responses and Wang et al. [40] integrate the search components into medical LLMs to help retrieve external knowledge for updating support evidence, which demonstrates domain retrieval is more important than domain pre-train. Nevertheless, there is a significant lack of effective methodologies to enhance the accuracy of LLMs in medical settings.
@@
Considering the language-specific knowledge that is highly relevant to local cultural, historical, political, and regional contexts, another crucial factor for the medical domain conversational search system is conducting multi/cross-lingual information retrieval [268, 269]. For example, traditional Chinese medicine, with its rich history deeply connected to cultural heritage, forms a unique medical search system [270, 143]. While a translation system can address some of these issues [41], it may prove ineffective due to the complexity of medical terminology and the challenges in achieving precise translations. Thus, exploring the fundamental mechanism of integrating language-specific medical knowledge and conversational search is promising and unexplored [42].
@@
To simulate the real-world scenario, Sharma et al. [43] build up a demo to show how financial agents can support stock market investors under conversational scenarios as shown in Figure 8 (b). Aim to develop sophisticated information systems for the financial domain and evaluate the model automatically, the follow-up studies focus on constructing the datasets based on conversational question answering tasks and storing the financial information in tabular [44, 45]. The key challenge lies in retrieving the relevant information from the structured data and generating the final answer by multi-hop reasoning, however, the performance of the models is still unsatisfactory because of data scarcity [46].
With huge pre-trained corpus, LLMs show superior capacity on reasoning ability but still fail in the financial domain, especially requiring retrieving numerical information [46]. To address such issues, [47] develop an LLM-based conversational search model for the financial domain, tailored for query intent classification and knowledge base labeling. However, the system has not been subjected to comprehensive information retrieval testing due to a lack of precise instructions and evaluation protocol. Xie et al. [48] further propose a comprehensive benchmark, instruction dataset, and LLMs for finance, facilitating the research field by providing systematic resources. Overall, the datasets, evaluation protocol, and model build-up technique constraint the development of financial systems to satisfy users’ special information needs, which require the domain-specific conversation search system to improve financial search and augment the response generation with correct retrieved information [262].
@@
Legal case retrieval is rapidly gaining prominence in the IR community due to its specific social impact, which aims to search for supporting evidence for a given query case. However, using a legal search system is challenging for those lacking domain knowledge because the user cannot fully express their specific information needs as illustrated in Figure 8 (c). To this end, Liu et al. [49] try to deploy a conversational search system into legal case retrieval and demonstrate its better effectiveness than the traditional ones. This is also attributed to the domain-specific linguistic phenomena, where the systematic patterns in language use, as explained by certain theories of conversational implicature, can sometimes explain why judges interpret legal texts in ways that differ from their literal meanings [50]. To further improve the performance of the system, a series of studies address different aspects, including developing agent-based action decision-making mechanisms [144, 51], re-rank long-context response [52], and learn to ask clarification question as introduced in Section 3.4.2.
@@
In addition to the aforementioned domains with text data only, the conversational search systems also facilitate the domains with different data forms. For example, the e-commercial domain involves various item attribution and user profile information, which requires dealing with product search, preference recommendation, and question answering based on item catalog matching, user modeling, and item-related knowledge understanding, respectively. Some recent studies [53, 145, 146] contribute to resource construction for the mixture search goals, entity retrieval, and scaling broad-coverage dialogue corpus, which enable the further exploration become much more feasible.
Besides, aiming to retrieve from multiple resources, Liao et al. [147] extend the conversational session to multi-modal scenarios, requiring the text and the image information for solving the downstream tasks. Li et al [148] emphasize more on the search task and further incorporate tabular evidence as retrieval sources. A typical example is shown in Figure 8 (d). The challenging task raises the difficulty of the alignment and reasoning among the retrieved evidence from different modalities, resulting in unsatisfactory performance by the previous state-of-the-art methods. To address it, MOQAGPT [149] is proposed based on a divide-and-conquer strategy to integrate the ranking results from various resources. More studies are required to facilitate this specific task setting.

### v2

@@ 4.4 Interpretation in Conversational Dense Retrieval
Although conversational dense retrieval has proven effective for conversational search, a significant limitation of these approaches is low interpretability, which impedes an intuitive understanding of model behaviors and hinders targeted improvements. To this end, the existing studies [255] attempt to unravel the behavior characteristics of various conversational dense retrieval models. For example, Mao et al. [154] propose LeCoRE, which enhances interpretability by creating a denoised and sparse lexical representation of the conversational session. Specifically, LeCoRE extends the SPLADE model [252], a lexical-based model to further transform latent representation into lexical representation in the vocabulary space, which provides a clear and understandable representation related to the search results. Besides, Christmann et al. [155] increase the model interpretability by combining information from multiple sources with user-comprehensible explanations for search results. Further, Cheng et al. propose CONVINV [156] to transform the opaque, high-dimensional session embeddings generated by conversational dense retrieval models into explicit, interpretable text that can be easily understood by humans. The transformation is achieved by a model trained to invert the session embeddings back into a textual form that is coherent and clear and maintains the retrieval performance of the original embeddings. Instead of explicitly conducting query reformulation for conversational search, more detailed investigations of the interpretability in dense retrieval are expected, which can avoid unfair or misleading search results.
@@ 4.5 Re-ranking in Conversational Search
In the context of re-ranking in conversational search, Kongyoung et al. [157] design monoQA to train a model with multi-task learning simultaneously for re-ranking retrieved passages and extracting answers via machine reading to improve both overall retrieval precision and answer accuracy. Ju et al. [30] attempt to improve re-ranking accuracy by employing a multi-view pseudo-labeling approach that integrates diverse contextual information. It utilizes a view-ensemble method to generate high-quality pseudo-labeled data for fine-tuning a conversational passage re-ranker. Hai et al. [140] design a re-ranking process after the conversational retrieval stage, which is enriched with keywords identified by the first-stage retriever. These keywords are expected to capture the helpful contextual information and re-rank the documents with useful conversational history. Owoicho et al. [158] investigate conversational search systems by providing feedback on system responses and answering clarifying questions, based on an initial description of the user’s information need. This work discusses the potential advancements in conversational search by developing more sophisticated feedback processing modules, e.g., conversational context-originated re-ranker.
@@ 4.6 Limitations and Discussion
Standard of Effective Pipeline. Though various pipelines with different mechanisms for conversational retrieval have been explored in the literature, how to establish a unified standard to construct and compare effective pipelines is still an open question, e.g., how to compare query reformulation approaches with conversational dense retrieval models. Besides, how to incorporate the multi-turn information, e.g., the historical retrieved results, into the current turn retrieval is also underexplored, as well as the necessity of deploying a re-ranker in the retrieval pipeline. These problems require the community to conduct further investigation on how to build a conversational retrieval pipeline with the associated historical context under a comparable standard. Besides, the rise of LLMs further complicates the pipeline standard. Generative retrieval models [256] or end-to-end RAG systems [257] inherently blend retrieval with context understanding and aspects of generation, making direct comparisons with traditional query reformulation or pipelined approaches difficult. Evaluating the retrieval component in such integrated systems requires assessing not just the relevance of retrieved passages but also their utility and impact on the final generated response [197].
@@
Explainability. Though some existing studies explore the explainability of conversational dense retrievers, the investigation of the interpretation studies is still limited. Thorough studies on conversational retrieval models can help the community understand the behavior of both the user and the retriever, which is essential for an interactive system. One potential solution is to conduct a comprehensive performance analysis that explores how different aspects of the conversation (e.g., topic switch and expansion utilization) affect retrieval effectiveness and user modeling.
@@ 5 Generation in Conversational Search
The development of generative models facilitates the evolution of conversational search systems that might enrich the user experience by producing contextually aware and personalized responses [3]. As the chat mode, i.e., multi-turn conversation, becomes a common practice when leveraging LLMs for content generation, the design of the new paradigms of the conversational search system by integrating the generation module brings much more attention. Based on powerful LLMs as generators, response generation is coupled with retrieval through Retrieval-Augmented Generation (RAG) paradigms. This integration means the generation component is not merely presenting information but actively synthesizing retrieved knowledge within the ongoing conversational context to meet the user’s information needs, further blurring the lines between traditional retrieval and generation stages.
@@
Conversational RAG [168] is not well-studied in the literature. The existing studies mainly focus on single-turn RAG, which only treats a single stand-alone query or converts the multi-turn context-dependent query by reformulation technique without leveraging historical information for RAG. In terms of the single-turn RAG paradigm, the generation component is pivotal for crafting responses that are both informed by external knowledge and coherent with the parametric ones. The retrieved documents usually served as external knowledge, and the generator model, usually the LLMs in existing studies, processes them with two key methodologies [259]: context curation [260, 261, 262, 263] and generator fine-tuning [264, 265], which try to adapt the consistency between IR and RAG [266, 267, 167] and satisfy the specific requirements of the domain or task, respectively. Besides, fine-tuning knowledge bases or enhancing LLMs’ internal knowledge could be the other alternatives to tune systems to meet specific conversational search expectations in generative paradigms. However, a single LLM, even with specific fine-tuning, could not be considered as a conversational search system without retrieval parts in this paper. This is because solely relying on a language model with parametric knowledge is not enough for information-seeking goals, and it is also the fundamental weakness of LLMs. Thus, the existing LLMs’ application requires integration with search engines (e.g., RAG), and there is no evidence showing that the functionality of the search component could be replaced.
@@
Regarding the generation for conversational search, the single-turn RAG systems overlook the historical information, e.g., historical search results, the context history, turn dependency, etc. Since conversational search systems must account for the ongoing dialogue and the evolving context of the conversation to produce responses that are not only relevant but also consistent with the user’s previous interactions, historical information might be essential to generate accurate, reliable, and personalized responses. Thus, it is unclear whether the additional conversational context process can further improve the generation results, even though some mechanisms with similar goals have already been applied in the retrieval procedure. Considering few studies are focusing on the concept of generation in conversational search, in this section, we aim to provide the potential research questions, guidance for unclear points, and corresponding solutions with different aspects to facilitate future research in this field. An overview of the conversational generation system is illustrated in Fig. 7.
@@
Figure 7: Overview of the conversational response generation system. To achieve more user-centric and reliable responses, it is worth further exploring, including historical context utilization, context dependency modeling, and knowledge attribution.
@@ 5.3 Conversational Knowledge Attribution
Although these methods enhance knowledge attribution and interpretability from different perspectives, they are not designed for conversational scenarios. Thus, it is unclear whether they can be applied to the conversational generation. For example, could these methods also be attributed to the historical evidence retrieved when injecting context information from previous turns? The potential exploration and improvement could follow the aspects below.
@@ 6 Domain-specific and User-centric Conversational Search
The rapid development of conversational search systems facilitates various applications in the general and diverse specific domains, including medical [274], finance [275], law [53], e-commercial [170], etc. The domain-specific systems might store the corresponding knowledge in different forms, e.g., tables, graphs, or multi-modal, which require the search systems to gather various information resources via reasoning skills. Aside from specific domain requirements, the system should cover the intended utilization of the conversational search model by real users. Thus, the designed system should also consider a user-centric perspective [200] in both dataset construction and evaluation criteria rather than only focusing on predefined search abilities. Besides, from the personalized perspective, since the backgrounds of users are different, the search results from the conversational systems should yield various results for the same query according to the user profiles with identified relevant information for each query turn [12]. In this section, we present the pioneering conversational search studies conducted in various specific domains and user-centric scenarios, and we conclude by offering a comprehensive discussion of the emerging limitations and challenges.
@@ 6.1 Domain-specific Conversational Search
Earlier studies [276, 277, 278] attempt to solve domain-specific issues via multi-domain/task learning based on task-originated conversation system. The key idea is to search cross-domain information for sharing by reasoning over the conversational history. Though these systems can somehow process various types of information across domains, they cannot achieve satisfactory performance when targeting each specific domain. Thus, various domain-specific models are necessarily proposed for each domain.
@@
The common practice of establishing a conversational search system in the medical domain is to extract relevant medical concepts from retrieved information within a conversational interface for response generation [35], which is illustrated in Fig. 8 (a). The medical domain presents unique challenges due to the significant consequences at stake, the need for extreme accuracy, the frequent use of complex terminology that is often unclear to non-experts, and privacy concerns. Nevertheless, existing conversational search is well-suited here because medical conditions often require discussing history, symptoms, and treatments over time, fitting the conversational paradigm. The variability in how medical concepts are expressed (i.e., the same concept might be written in numerous non-canonical forms) poses challenges for accurately searching texts with relevant concepts and for modeling conversational context, particularly with co-reference and omission [36]. Thus, effective systems must bridge the gap between patient language and clinical terminology.
@@
LLM-based approaches. With the emergence of LLMs, the medical domain conversational search systems paradigm encounters new opportunities and challenges [279, 274]. The available medical corpus [37, 38] enables continual training of the existing LLMs to enhance their superb capability in the associated domain. For example, Toma et al. [39] propose an open-source expert-level medical language model by encoding the knowledge based on conversational context. Bao et al. [40] construct high-quality supervised fine-tuning datasets based on dialogue and knowledge graphs to bridge LLMs with real-world medical consultation. However, the generative LLMs remain an unyielding concern in returning accurate responses, especially when the stored knowledge is out-to-date. To mitigate such issues, Pal et al. [41] propose medical domain hallucination detection approaches for LLMs to reduce inaccurate responses and Wang et al. [42] integrate the search components into medical LLMs to help retrieve external knowledge for updating support evidence, which demonstrates domain retrieval is more important than domain pre-train. Although integrating external knowledge is promising, the current LLM-based approaches still struggle with hallucination issues, which is especially dangerous in medicine. Furthermore, existing medical domain benchmarks often focus on standardized question-answering tasks rather than simulating realistic, nuanced patient consultations. Evaluating safety, reliability, and the ability to handle sensitive information appropriately remains critical and underdeveloped [280]. Thus, there is a significant lack of effective methodologies to enhance the LLMs to achieve users’ specific goals in medical settings.
@@
Language-specific knowledge in Medical Domain. Considering the language-specific knowledge that is highly relevant to local cultural, historical, political, and regional contexts, another crucial factor for the medical domain conversational search system is conducting multi/cross-lingual information retrieval [281, 282]. For example, with its rich history deeply connected to cultural heritage, forms a unique medical search system [283, 169]. While a translation system can address some of these issues [43], it may prove ineffective due to the complexity of medical terminology and the challenges in achieving precise translations. Integrating diverse medical knowledge, e.g., for Traditional Chinese Medicine, requires more than translation, which demands culturally aware models capable of understanding different diagnostic paradigms. Thus, exploring the fundamental mechanism of integrating language-specific medical knowledge and conversational search is promising and unexplored [44].
@@
Finance with numerical data demands high precision and often involves complex reasoning over structured and unstructured information, such as tables and reports [284, 285]. The information needs of the users might be real-time data or performing transactional tasks. Conversational search systems can be suitable for guiding users through complex financial products or explaining market trends, but must handle numerical reasoning robustly and ensure information timeliness [48]. A series of studies have been conducted in the financial domain in the context of conversational search. To simulate the real-world scenario, Sharma et al. [45] build up a demo to show how financial agents can support stock market investors under conversational scenarios as shown in Fig. 8 (b). Aim to develop sophisticated information systems and evaluate the model automatically, the follow-up studies focus on constructing the datasets based on conversational question answering tasks and storing the financial information in tabular [46, 47]. The key challenge lies in retrieving the relevant information from the structured data and generating the final answer by multi-hop reasoning, however, the performance of the models is still unsatisfactory because of data scarcity [48].
With a huge pre-trained corpus, LLMs show superior capacity in reasoning ability but still fail in the financial domain, especially requiring retrieving numerical information [48]. To address such issues, [49] develop an LLM-based conversational search model for the financial domain, tailored for query intent classification and knowledge base labeling. However, the system has not been subjected to comprehensive information retrieval testing due to a lack of precise instructions and evaluation protocol. Xie et al. [50] further propose a comprehensive benchmark, instruction dataset, and LLMs for finance, facilitating the research field by providing systematic resources. Overall, current systems show promise but often struggle with multi-step numerical reasoning and integrating real-time data feeds. The existing datasets, benchmark, and evaluation protocol are valuable but may not fully capture the complexity and dynamism of real-world financial queries. These factors constrain the development of financial systems to satisfy users’ special information needs [275]. Moreover, ensuring that the generated advice is both accurate and compliant with financial regulations represents a significant challenge that has been largely neglected in existing research evaluations.
@@
Legal case retrieval is rapidly gaining prominence in the IR community due to its specific social impact, which aims to search for supporting evidence for a given query case. However, using a legal search system is challenging for those lacking domain knowledge because the user cannot fully express their specific information needs as illustrated in Fig. 8 (c). Specifically, the legal domain requires retrieving precise supporting evidence, e.g., statutes, case law, which often involves complex logical reasoning and the interpretation of highly technical, terminology-rich texts. The ambiguity in law cases is common, while interpretations have significant impacts. To this end, Liu et al. [51] try to deploy a conversational search system into legal case retrieval and demonstrate its better effectiveness than the traditional ones. This is also attributed to the domain-specific linguistic phenomena, where the systematic patterns in language use, as explained by certain theories of conversational implicature, can sometimes explain why judges interpret legal texts in ways that differ from their literal meanings [52]. To further improve the performance of the system, a series of studies address different aspects, including developing agent-based action decision-making mechanisms [124, 53], re-rank long-context response [54], and learn to ask clarification questions as introduced in Section 3.4.2. While achieving some improvements, current conversational search in the legal domain faces challenges in ensuring the retrieved evidence is contextually appropriate and legally sound. The ability to handle long, complex legal documents effectively within a conversational context [54] is also a key area for improvement.
@@
In addition to the aforementioned domains with text data only, the conversational search systems also facilitate domains with different data forms. For example, the e-commercial domain involves various item attribution and user profile information, which requires dealing with product search, preference recommendation, and question answering based on item catalog matching, user modeling, and item-related knowledge understanding, respectively. Some recent studies [55, 170, 171] contribute to resource construction for the mixture search goals, entity retrieval, and scaling a broad-coverage dialogue corpus, which enables further exploration to become much more feasible.
Besides, aiming to retrieve from multiple resources, Liao et al. [172] extend the conversational session to multi-modal scenarios, requiring the text and the image information for solving the downstream tasks. Li et al [173] emphasize more on the search task and further incorporate tabular evidence as retrieval sources. A typical example is shown in Fig. 8 (d). The challenging task raises the difficulty of the alignment and reasoning among the retrieved evidence from different modalities, resulting in unsatisfactory performance by the previous state-of-the-art methods. To address it, MOQAGPT [174] is proposed based on a divide-and-conquer strategy to integrate the ranking results from various resources. More studies are required to facilitate this specific task setting.